# Rolling Diffusion Models

### v1

@@ 1 Introduction
After these successes in these domains, interest in developing diffusion models for time sequences has grown. Prominent recent large-scale works include, e.g., Imagen Video (Ho et al., 2022a), Stable Diffusion Video (StabilityAI, 2023). Other impressive results for generating video data have been achieved, e.g., by (Blattmann et al., 2023; Ge et al., 2023; Harvey et al., 2022; Singer et al., 2022; Ho et al., 2022b). Applications of sequential generative modeling outside video include, e.g., fluid mechanics or weather and climate modeling (Price et al., 2023; Meng et al., 2022; Lippe et al., 2023).
What’s common across many of these works is that they treat the temporal axis as an “extra spatial dimension”. That is, they treat the video as a 3D tensor of shape $K\times H\times W$. This has several downsides. First, the memory and computational requirements can quickly grow infeasible if one wants to generate long sequences. Second, one is typically interested in being able to roll out generation for a variable number of time steps. Therefore, an alternative angle is a fully autoregressive approach by conditioning on a sequence of input frames and simulating a single output frame, which is then concatenated to the input frames, upon which the recursion can continue. In this case, one has to traverse the entire denoising diffusion chain for every single frame, which is computationally intensive. Additionally, iteratively sampling single frames leads to quick autoregressive error accumulation. A middle ground can be found by jointly generating blocks of frames. However, in this block-autoregressive case, a diffusion model would use the same number of denoising steps for every frame. This is suboptimal since, given a sequence of input frames, the uncertainty about the first upcoming few is much lower than the later ones. Finally, both methods sample frames only jointly with earlier frames, which is potentially a suboptimal parameterization.
In this work, we propose a new framework called Rolling Diffusion, a method that explicitly corrupts data from past to future. This is achieved by reparameterizing the global diffusion time to a local time for each frame. It turns out that by doing this, one can (apart from boundary conditions) completely focus on a local sliding window sequential denoising process. This has several temporal inductive biases, alleviating some of the abovementioned issues.
@@
    The model only has to predict the low frequencies (corresponding to their global visual features) for distant frames, whereas the high-frequency details are generated for temporally near frames.
@@
These merits are empirically demonstrated in, among others, a video prediction experiment using the Kinetics-600 video dataset and in an experiment involving chaotic fluid mechanics simulations.
@@ 2.1 Diffusion
Diffusion models consist of a process that destroys data stochastically, named the diffusion process and a generative process called the denoising process. Let ${\bm{z}}_{t}\in{\mathbb{R}}^{D}$ denote a latent variable over a diffusion dimension $t\in[0,1]$. Throughout this work, we will refer to $t$ as the global (diffusion) time. Given a datapoint ${\bm{x}}\in{\mathbb{R}}^{D}$, ${\bm{x}}\sim q({\bm{x}})$, the diffusion process is designed so that ${\bm{z}}_{0}\approx{\bm{x}}$ and ${\bm{z}}_{1}\sim\mathcal{N}(0,1)$ via the distribution:
@@
$\operatorname{SNR}(t):=\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}}\,,$ (2)
@@
Given the noising process, it can be shown (Sohl-Dickstein et al., 2015) that the true (i.e., optimal) denoising distribution for a single datapoint ${\bm{x}}$ from time $t$ to time $s$ is given by
@@
The diffusion objective can be expressed as a KL-divergence between the diffusion process and the denoising process, i.e. $\mathrm{KL}(q({\bm{x}},{\bm{z}}_{0},\ldots,{\bm{z}}_{1})|p({\bm{x}},{\bm{z}}_{0},\ldots,{\bm{z}}_{1}))$ which simplifies to (Kingma et al., 2021):
$\mathcal{L}_{\theta}({\bm{x}})$ $:=\mathbb{E}_{t\sim U(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,1)}\left[w(t)||{\bm{x}}-f_{\theta}({\bm{z}}_{t,{\bm{{\epsilon}}}},t)||^{2}\right]$
@@
where $\mathcal{L}_{\mathrm{prior}}$ and $\mathcal{L}_{\mathrm{data}}$ are typically negligible. The weighting $w(t)$ can be freely specified. In practice, it was found that specific weightings of the loss result in better sample quality (Ho et al., 2020). This is the case for, e.g., ${\bm{{\epsilon}}}$-loss (Ho et al., 2020), which corresponds to $w(t)=\mathrm{SNR}(t)$.
@@ 2.2 Diffusion for temporal data
If one is interested in generation of temporal data, potentially indefinitely (or beyond typical hardware constraints), one must consider (autoregressive) conditional extension of previously generated data. I.e., given an initial sample ${\bm{x}}^{k}\sim q({\bm{x}})$ at a temporal index $k$, we want to estimate and sample the conditional distribution $p({\bm{x}}^{k+1}|{\bm{x}}^{k})$. This process can then be extended to videos of arbitrary lengths. As discussed in Section 1, it is not yet clear what kinds of parameterization choices are optimal to estimate this conditional distribution. Further, no temporal inductive bias is typically baked into the denoising process.
@@ 3 Rolling Diffusion Models
Figure 1: Left: an illustration of a global rolling diffusion process and its local time reparameterization. The global diffusion time $t$ (vertical axis) is mapped to a local time $t_{k}$ for a frame $k$ (horizontal axis). The local time is then used to compute the diffusion parameters $\alpha_{t_{k}}$ and $\sigma_{t_{k}}$. On the right, we show how the same local schedule can be applied to each sequence of frames based on the frame index $w$. The nontrivial part of sampling the generative process only occurs in the sliding window as it gets shifted over the sequence.
@@ 3.1 A global perspective
Let ${\bm{x}}\in\mathbb{R}^{D\times K}$ be a time series datapoint where $K$ denotes the number of frames and $D$ the dimensionality of each frame. The core idea that allows rolling diffusion is a reparameterization of the diffusion time $t$ to a frame-dependent local (frame-dependent) time: i.e.,
@@
Note that we still require $t_{k}\in[0,1]$ for all ${k\in\{0,\dots,K-1\}}$. Furthermore, we still have a monotonically decreasing signal-to-noise schedule, ensuring a well-defined diffusion process. However, we now (effectively) have a different signal-to-noise schedule for each frame. In this work, we also always have $t_{k}\leq t_{k+1}$ i.e., the local time of a frame is smaller than the local time of the next frame. This means we add more noise to future frames: a natural temporal inductive bias. Note that this is not strictly required; one could also have a reverse-time inductive bias or a mixture. An example of such a reparameterization is shown in Figure 1 (left). We depict a map that takes a global diffusion time $t$ (vertical axis) and a frame index $k$ (horizontal axis), and computes a local time $t_{k}$, indicated with a color intensity.
@@
This categorization can be motivated using the schedule depicted in Figure 1. Given, for example, $t=0.5$ and $s=0.375$, we see that the first frame $k=0$ falls in the first category. At this point in time, ${\bm{z}}_{t_{0}}={\bm{z}}_{s_{0}}$ are identical given that $\lim_{t\to 0^{+}}\log\mathrm{SNR}(t)=\infty$. On the other hand, the last frame $k=K-1$ (31 in the figure) falls in the second category, i.e., both ${\bm{z}}_{t_{K-1}}$ and ${\bm{z}}_{s_{K-1}}$ are distributed as independent standard Gaussians, given that $\lim_{t\to 1^{-}}\log\mathrm{SNR}(t)=-\infty$. Finally, the frame $k=16$ falls in the third, most interesting category: the sliding window. As such, observe that the true denoising process can be factorized as:
@@
$p_{\theta}({\bm{z}}_{s}^{\text{win}}|{\bm{z}}_{t})$ $=p_{\theta}({\bm{z}}_{s}^{\text{win}}|{\bm{z}}^{\text{clean}}_{t},{\bm{z}}^{\text{win}}_{t})\,.$ (16)
In practice, we approximate $p_{\theta}({\bm{z}}_{s}^{\text{win}}|{\bm{z}}^{\text{clean}}_{t},{\bm{z}}^{\text{win}}_{t}):\approx p_{\theta}({\bm{z}}_{s}^{\text{win}}|{\bm{z}}^{\widehat{\text{clean}}},{\bm{z}}^{\text{win}}_{t})$, where ${\bm{z}}^{\widehat{\text{clean}}}$ denotes a specific (potentially empty) subset of ${\bm{z}}^{\text{clean}}_{t}$ used for extra conditioning. This typically includes a few frames slightly before the current sliding window.
In Appendix D, we formalize the above and show that the objective results in
@@
$L_{\text{win},\theta}:=\sum_{k\in\text{win}(t)}a(t_{k})||{\bm{x}}^{k}-f^{k}_{\theta}({\bm{z}}_{t,{\bm{{\epsilon}}}}^{\text{win}},{\bm{z}}_{t,{\bm{{\epsilon}}}}^{\widehat{\text{clean}}},t)||^{2}\,,$
where we suppress some arguments for notational convenience.
Observe Figure 1 again. After training is completed, we can essentially sample from the generative model by traversing the image with the sliding window from the top left to the bottom right. This allows us to completely focus on the local environment, as discussed in the next section.
@@ 3.2 A local perspective
In the last section, we saw that rolling diffusion allows us to focus the generative process as well as training solely on frames that are in the sliding window. As such, we now have $t_{\text{win}}(t):[0,1]\to[0,1]$ denote a time reparameterization subject to the usual constraints. Note, though, that the shared time-component $t$ is now defined locally. Specifically, running the denoising chain from $t=1$ to $t=0$ will only sample a rolling window such that the first frame is completely noiseless and the following frames still contain some noise. In contrast, the global process described earlier denoises an entire video.
In principle, the design space of pure rolling diffusion models is enormous. For this reason, we make the following sensible assumptions in addition to the earlier constraints on the signal-to-noise schedule to allow for a sliding window sampling procedure. We would like $t_{\text{win}}$ to be:
@@
    consistent under moving the window, meaning that $t_{\text{win}}(1)$ should equal $t_{\text{win}}(0)$ of the next frame.
Let $W<K$ be the size of the sliding window, and $w\in\{0,\dots,W-1\}$ be the local indices of the frames. To satisfy the first assumption, we define the schedule in terms of the local index $w$ (see Figure 1 (right)).
$t_{\text{win}}:=t^{W}_{w},$ (18)
where $t_{w}^{W}:[0,1]\to[0,1]$ is a monotonically increasing function. For the second, we know $t^{W}_{w}$ must have
$t^{W}_{w}=g((w+t)/W),$ (19)
@@
$t_{w}^{\text{lin}}=(w+t)/W\,.$ (20)
where $w\in\{0,\dots,W-1\}$. See Figure 1 (right) for an illustration of how this local schedule is applied to each sequence of frames. Observe that
$t_{w}^{\text{lin}}\in[w/W,(w+1)/W]\subseteq[0,1]\,.$ (21)
One can extend the linear local time to include clean conditioning frames. Let $n_{\text{cln}}$ denote the number of clean frames (chosen as a hyperparameters), then the local time for a frame $w$ is:
@@
where $\operatorname{clip}:\mathbb{R}\to[0,1]$ clips value between 0 and 1. Although it may seem that this contradicts Equation 19, this is a simple reparametrization.

### v2

@@ 1 Introduction
After these successes in these domains, interest in developing diffusion models for time sequences has grown. Prominent recent large-scale works include, e.g., Imagen Video (Ho et al., 2022a), Stable Diffusion Video (StabilityAI, 2023). Other impressive results for generating video data have been achieved by, e.g., (Blattmann et al., 2023; Ge et al., 2023; Harvey et al., 2022; Singer et al., 2022; Ho et al., 2022b). Applications of sequential generative modeling outside video include, e.g., fluid mechanics or weather and climate modeling (Price et al., 2023; Meng et al., 2022; Lippe et al., 2023).
What is common across many of these works is that they treat the temporal axis as an ‘extra spatial dimension’. That is, they treat the video as a 3D tensor of shape $K\times H\times W$. This has several downsides. First, the memory and computational requirements can quickly grow infeasible if one wants to generate long sequences. Second, one is typically interested in being able to roll out the sampling process for a variable number of time steps. Therefore, an alternative angle is a fully autoregressive approach by conditioning on a sequence of input frames and simulating a single output frame, which is then concatenated to the input frames, upon which the recursion can continue. In this case, one has to traverse the entire denoising diffusion sampling chain for every single frame, which is computationally intensive. Additionally, iteratively sampling single frames leads to quick autoregressive error accumulation. A middle ground can be found by jointly generating blocks of frames. However, in this block-autoregressive case, a diffusion model would use the same number of denoising steps for every frame. This is suboptimal since, given a sequence of conditioning frames, the generative uncertainty about the next few is much lower than the frames further into the future. Finally, both methods sample frames only jointly with earlier frames, which is potentially a suboptimal parameterization.
Figure 1: Overview of the Rolling Diffusion rollout sampling procedure. The input to the model contains some conditioning and a sequence of partially denoised frames. The model then denoises the frames by a small amount. After denoising, the sliding window shifts, and the fully denoised frames are concatenated with the conditioning. This process is repeats until the desired number of frames is generated. Example video taken from the Kinetics-600 dataset (Kay et al., 2017) (CC BY 4.0).
In this paper, we propose a new framework called Rolling Diffusion, a method that explicitly corrupts data from past to future. This is achieved by reparameterizing the global diffusion time to a local time for each frame. It turns out that by doing this, one can (apart from boundary conditions) completely focus on a local sliding window sequential denoising process. This has several temporal inductive biases, alleviating some of the abovementioned issues.
@@
    In denoising diffusion models, the model output tends to contain low-frequency information in high noise regimes and includes high-frequency information only when corruptions are light. In our framework, the noise strength is higher for frames that are further from the conditioning. As such, the model only needs to predict low-frequency information (i.e., global structures) for frames further into the future; high-frequency information gets included as frames move closer to the present.
@@
These merits are empirically demonstrated in, among others, a video prediction experiment using the Kinetics-600 video dataset (Kay et al., 2017) and in an experiment involving chaotic fluid mechanics simulations.
Figure 2: Left: an illustration of a global rolling diffusion process and its local time reparameterization. The global diffusion denoising time $t$ (vertical axis) is mapped to a local time $t_{k}$ for a frame $k$ (horizontal axis). The local time is then used to compute the diffusion parameters $\alpha_{t_{k}}$ and $\sigma_{t_{k}}$. On the right, we show how the same local schedule can be applied to each sequence of frames based on the frame index $w$. The nontrivial part of sampling the generative process only occurs in the sliding window as it gets shifted over the sequence.
@@ 2.1 Diffusion
Diffusion models consist of a process that destroys data stochastically, named the ‘diffusion process’, and a generative process called the denoising process. Let ${\bm{z}}_{t}\in{\mathbb{R}}^{D}$ denote a latent variable over a diffusion dimension $t\in[0,1]$. We refer to $t$ as the global (diffusion) time, which will determine the amount of noise added to the data. Given a datapoint ${\bm{x}}\in{\mathbb{R}}^{D}$, ${\bm{x}}\sim q({\bm{x}})$, the diffusion process is designed so that ${\bm{z}}_{0}\approx{\bm{x}}$ and ${\bm{z}}_{1}\sim\mathcal{N}(0,1)$ via the distribution:
@@
$\operatorname{SNR}(t):=\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}}$ (2)
@@
Given the noising process, it can be shown (Sohl-Dickstein et al., 2015) that the true (i.e., optimal) denoising distribution for a single datapoint ${\bm{x}}$ from time $t$ to time $s$ ($s\leq t)$ is given by
@@
The diffusion objective can be expressed as a KL-divergence between the diffusion process and the denoising process, i.e. $D_{\mathrm{KL}}(q({\bm{x}},{\bm{z}}_{0},\ldots,{\bm{z}}_{1})\,||\,p({\bm{x}},{\bm{z}}_{0},\ldots,{\bm{z}}_{1}))$ which simplifies to (Kingma et al., 2021):
$\mathcal{L}_{\theta}({\bm{x}})$ $:=\mathbb{E}_{t\sim U(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,1)}\left[a(t)||{\bm{x}}-f_{\theta}({\bm{z}}_{t,{\bm{{\epsilon}}}},t)||^{2}\right]$
@@
where $\mathcal{L}_{\mathrm{prior}}$ and $\mathcal{L}_{\mathrm{data}}$ are typically negligible. The weighting $a(t)$ can be freely specified. In practice, it was found that specific weightings of the loss result in better sample quality (Ho et al., 2020). This is the case for, e.g., $\epsilon$-loss, which corresponds to $a(t)=\mathrm{SNR}(t)$.
@@ 2.2 Diffusion for temporal data
If one is interested in generation of temporal data beyond typical hardware constraints, one must consider (autoregressive) conditional extension of previously generated data. I.e., given an initial sample ${\bm{x}}^{k}$ at a temporal index $k$, we want to sample a (faithful) conditional distribution $p({\bm{x}}^{k+1}|{\bm{x}}^{k})$. This process can then be extended to videos of arbitrary lengths. As discussed in Section 1, it is not yet clear what kinds of parameterization choices are optimal to estimate this conditional distribution. Further, no temporal inductive bias is typically baked into the denoising process.
@@ 3.1 A global perspective
Let ${\bm{x}}\in\mathbb{R}^{D\times K}$ be a time series datapoint where $K$ denotes the number of frames and $D$ the dimensionality of each frame. The core idea that allows rolling diffusion is a reparameterization of the diffusion (denoising) time $t$ to a frame-dependent local (frame-dependent) time: i.e.,
@@
Note that we still require $t_{k}\in[0,1]$ for all ${k\in\{0,\dots,K-1\}}$. Furthermore, we still have a monotonically decreasing signal-to-noise schedule, ensuring a well-defined diffusion process. However, we now have a different signal-to-noise schedule for each frame. In this work, we also always have $t_{k}\leq t_{k+1}$, i.e., the local denoising time of a given frame is smaller than the local time of the next frame. This means we add more noise to future frames: a natural temporal inductive bias. Note that this is not strictly required; one could also have a reverse-time inductive bias or a mixture. An example of such a reparameterization is shown in Figure 2 (left). We depict a map that takes a global diffusion time $t$ (vertical axis) and a frame index $k$ (horizontal axis), and computes a local time $t_{k}$, indicated with a color intensity.
@@
Note that here, $s$ and $t$ are both diffusion time-steps (corresponding to certain SNR levels), while $k$ denotes a frame index. This categorization can be motivated using the schedule depicted in Figure 2. Given, for example, $t=0.5$ and $s=0.375$, we see that the first frame $k=0$ falls in the first category. At this point in time, ${\bm{z}}_{t_{0}}={\bm{z}}_{s_{0}}$ are identical given that $\lim_{t\to 0^{+}}\log\mathrm{SNR}(t)=\infty$. On the other hand, the last frame $k=K-1$ (31 in the figure) falls in the second category, i.e., both ${\bm{z}}_{t_{K-1}}$ and ${\bm{z}}_{s_{K-1}}$ are distributed as independent standard Gaussians, given that $\lim_{t\to 1^{-}}\log\mathrm{SNR}(t)=-\infty$. Finally, the frame $k=16$ falls in the third, most interesting category: the sliding window. As such, observe that the true denoising process can be factorized as:
@@
$p_{\theta}({\bm{z}}_{s}^{\text{win}}|{\bm{z}}_{t})$ $=p_{\theta}({\bm{z}}_{s}^{\text{win}}|{\bm{z}}^{\text{clean}}_{t},{\bm{z}}^{\text{win}}_{t})$ (16)
$:\approx p_{\theta}({\bm{z}}_{s}^{\text{win}}|\widehat{{\bm{z}}_{t}^{\text{clean}}},{\bm{z}}^{\text{win}}_{t})$ (17)
where $\widehat{{\bm{z}}_{t}^{\text{clean}}}$ denotes a specific subset of ${\bm{z}}^{\text{clean}}_{t}$, typically including a few frames slightly before the current sliding window.
In Appendix A, we motivate, in addition to the arguments above, the following loss function:
@@
$L_{\text{win},\theta}:=\sum_{k\in\text{win}(t)}a(t_{k})||{\bm{x}}^{k}-f^{k}_{\theta}({\bm{z}}_{t,{\bm{{\epsilon}}}}^{\text{win}},{\bm{z}}_{t,{\bm{{\epsilon}}}}^{\text{clean}},t)||^{2}\,,$
where we suppress some arguments for notational convenience. Here, ${\bm{z}}_{t,{\bm{{\epsilon}}}}$ denotes a noised version of ${\bm{x}}$ as a function of $t$ and ${\bm{{\epsilon}}}$ and $a(t_{k})$ is a weighting function leading to, e.g., the usual ‘simple’ $\epsilon$-MSE loss, $v$-MSE loss, or $x$-MSE loss.
Observe Figure 2 again. After training is completed, we can essentially sample from the generative model by traversing the image with the sliding window from the top left to the bottom right.
@@ 3.2 A local perspective
In the previous section, we discussed how rolling diffusion enables us to concentrate entirely on frames within a sliding window. Instead of using $t$ to represent the global diffusion time, which determines the noise level for all frames, we now redefine $t$ to determine the noise level for each frame in a smaller subsequence. Specifically, running the denoising chain from $t=1$ to $t=0$ will sample a sequence such that the first frame is completely denoised, but the subsequent frames still retain some noise. In contrast, the global process described earlier denoises an entire video.
Similar to before, we reparameterize $t$ to allow for different noise levels for each frame in the sliding window. This reparameterization should be:
@@
    consistent under moving the window, meaning that the noise level for the current frame $w$ when $t=0$ should match the noise level at $w-1$ at $t=1$. This consistency enables seamless denoising as the window slides, ensuring that each frame is progressively denoised while shifting positions.
Let $W<K$ be the size of the sliding window, and $w\in\{0,\dots,W-1\}$ be the local indices of the frames. To satisfy the first assumption, we define the schedule in terms of the local index $w$ (see Figure 2 (right)):
$t\mapsto t^{W}_{w}\,.$ (19)
For the second, we know $t^{W}_{w}$ must have
$t^{W}_{w}=g\left(\frac{w+t}{W}\right)$ (20)
@@
$t_{w}^{\text{lin}}=\frac{w+t}{W}\,.$ (21)
See Figure 2 (right) for an illustration of how this local schedule is applied to each sequence of frames. Observe that $t_{w}^{\text{lin}}\in[w/W,(w+1)/W]\subseteq[0,1]$. As such, we can directly use our SNR schedule to compute the diffusion parameters for each frame.
One can extend the linear local time to include clean conditioning frames. Let $n_{\text{cln}}$ denote the number of clean frames (chosen as a hyperparameter), then the local time for a frame $w$ is:
@@
where $\operatorname{clip}:\mathbb{R}\to[0,1]$ clips value between 0 and 1.
3.3 Boundary conditions
While framing rolling diffusion purely from a local perspective is convenient for training and sampling, it introduces some complicated edge conditions. Given the linear local time reparameterization $t^{\text{lin}}_{w}$, we have that given the diffusion time $t$ running from $1$ to 00, the local times run from $(\frac{1}{W},\frac{2}{W},\ldots\frac{W}{W})$ to $(\frac{0}{W},\frac{1}{W},\ldots,\frac{W-1}{W})$. This means that using this setting, the signal-to-noise ratios are never minimal for all frames, meaning we cannot start sampling from a completely noisy state. Visually, in Figure 2, the rolling sampling procedure can be seen as moving the sliding window over the diagonal from top left to bottom right such that the local times of the frames remain invariant upon shifting. However, this means that placing the window at the very left edge still results in having partially denoised frames.
To account for this, we co-train the rolling diffusion model with an additional schedule that can handle this boundary condition:
$t_{w}^{\text{init}}:=\operatorname{clip}\left(\frac{w}{W}+t\right)\,.$ (23)
This init noise schedule can start from initial random noise and generates a video in the ‘rolling state’. That is, at diffusion time $t=1$, this will put all frames to maximum noise, and at $t=0$ the frames will be in the rolling state. To be precise, it starts from local times $(1,1,\ldots 1)$ and denoises to $(0,\frac{1}{W},\frac{2}{W},\dots,\frac{W-1}{W})$, after which we can start using the previously described $t_{w}^{\text{lin}}$ schedule. From a visual perspective, in Figure 2, this corresponds to placing the window at the upper left corner and moving it down vertically, until it reaches the stage where it can start moving diagonally.
On the interval $[0,\frac{1}{W}]$, this schedule contains the previous local schedule $t_{w}^{\text{lin}}$ as a special case. To see this, note that $t_{w}^{\text{init}}(\frac{1}{W})=\left(\frac{1}{W},\frac{2}{W},\dots,1\right)$, which is the same as $t_{w}^{\text{lin}}(1)$. Similarly, one can check the case for $t_{w}^{\text{init}}(0)$. This means that the model could be trained solely with $t_{w}^{\text{init}}$ to handle the boundaries as well as being able to roll out indefinitely. At sampling time, one then still uses the $t_{w}^{\text{init}}$ schedule but restricted to $[0,\frac{1}{W}]$. The caveat, however, is that the $t^{\text{lin}}_{w}$ schedule only gets selected $1/W$ of the time during training (assuming $t\sim U(0,1)$). In contrast, this schedule is used almost exclusively at test time, with the exception being at the boundary. As such, we find it beneficial to oversample the $t_{w}^{\text{lin}}$ schedule during training based on a Bernoulli rate $\beta$.